If f(x) = 4x - 6, then f(3) = _____?
Answer by rapaljer(4671)

f(x)=4x-6

f(3) means to plug in the value of 3 in place of the x.
f(3)=4*3-6
f(3)=12-6=6

I do have a page on my website that might be a better explanation for you than the one you have been trying to understand. To get to my website, click on my tutor name "rapaljer" anywhere in algebra.com. There you will see a link to my website. From my Homepage, look for "Basic, Intermediate, and College Algebra." Choose "Basic Algebra", and look for "Chapter 4," "Section 4.09 Functional Notation." This is my own non-traditional explanation of a topic that is not nearly as difficult as it may seem. I believe math can be very EASY when it is taught right!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus